# Researcher Population Pyramids for Tracking Global Demographic and Gender Trajectories

### v1

@@ Acknowledgments
Country AUC F1 GAR AUC F1 GAR
Albania 0.984 0.963 0.907 0.950 0.963 0.861
Algeria 0.965 0.951 0.906 0.945 0.953 0.879
Australia 0.995 0.978 0.956 0.992 0.994 0.911
Austria 0.996 0.984 0.971 0.995 0.996 0.947
Bahrain 0.998 0.997 0.959 0.969 0.969 0.858
Belarus 0.999 0.999 0.967 0.997 0.998 0.978
Belgium 0.982 0.971 0.880 0.979 0.985 0.880
Bosnia and Herzegovina 0.986 0.974 0.908 0.979 0.978 0.908
Bulgaria 0.998 0.993 0.990 0.987 0.991 0.866
Canada 0.987 0.973 0.906 0.984 0.986 0.884
Cyprus 0.976 0.974 0.819 0.971 0.970 0.889
Czechia 0.999 0.995 0.959 0.999 0.999 0.931
Denmark 0.989 0.977 0.908 0.987 0.987 0.879
Egypt 0.988 0.968 0.939 0.945 0.944 0.791
Estonia 0.998 0.994 0.962 0.991 0.992 0.849
France 0.995 0.986 0.938 0.991 0.995 0.880
Germany 0.995 0.991 0.939 0.995 0.995 0.936
Ghana 0.994 0.982 0.929 0.975 0.986 0.886
Iceland 0.999 0.997 0.986 0.988 0.990 0.910
Iran, Islamic Republic of 0.981 0.974 0.852 0.933 0.941 0.812
Iraq 0.836 0.909 0.426 0.885 0.913 0.690
Ireland 0.999 0.997 0.964 0.998 0.998 0.943
Israel 0.965 0.953 0.856 0.941 0.952 0.816
Italy 0.999 0.999 0.999 0.999 0.999 0.926
Jamaica 0.978 0.969 0.836 0.964 0.972 0.789
Japan 0.981 0.956 0.934 0.967 0.970 0.808
Jordan 0.995 0.985 0.967 0.951 0.954 0.801
Kenya 0.914 0.895 0.920 0.907 0.896 0.933
Kuwait 0.985 0.968 0.912 0.928 0.942 0.753
Lebanon 0.997 0.992 0.959 0.926 0.931 0.866
Lesotho 0.975 0.965 0.867 0.687 0.867 0.276
Lithuania 0.994 0.991 0.996 0.994 0.994 0.946
Moldova, Republic of 0.993 0.986 0.973 0.985 0.986 0.941
Montenegro 0.983 0.955 0.918 0.955 0.959 0.877
Morocco 0.989 0.985 0.896 0.973 0.976 0.912
Netherlands 0.981 0.973 0.898 0.980 0.984 0.880
New Zealand 0.998 0.992 0.952 0.989 0.991 0.896
Nigeria 0.954 0.923 0.866 0.933 0.952 0.673
Norway 0.991 0.980 0.983 0.984 0.986 0.890
Oman 0.999 0.997 0.969 0.908 0.918 0.797
Papua New Guinea 0.983 0.969 0.897 0.946 0.970 0.774
Philippines 0.981 0.978 0.861 0.958 0.968 0.803
Poland 0.992 0.990 0.978 0.991 0.993 0.951
Portugal 0.996 0.992 0.923 0.992 0.994 0.924
Romania 0.980 0.967 0.921 0.962 0.973 0.819
Russian Federation 0.999 0.996 0.996 0.998 0.999 0.947
Saudi Arabia 0.981 0.972 0.906 0.878 0.901 0.767
Serbia 0.997 0.994 0.974 0.984 0.979 0.943
Somalia 0.851 0.798 0.816 0.965 0.970 0.818
South Africa 0.979 0.962 0.871 0.919 0.966 0.648
Spain 0.990 0.978 0.942 0.986 0.984 0.922
Sri Lanka 0.884 0.927 0.487 0.858 0.930 0.572
Sweden 0.996 0.988 0.936 0.988 0.992 0.921
Switzerland 0.999 0.992 0.990 0.999 0.999 0.948
Syrian Arab Republic 0.913 0.913 0.751 0.941 0.951 0.769
Tunisia 0.965 0.953 0.871 0.964 0.966 0.839
Turkey 0.990 0.975 0.930 0.972 0.977 0.854
United Arab Emirates 0.995 0.987 0.957 0.942 0.950 0.822
United Kingdom 0.998 0.991 0.966 0.996 0.998 0.897
United States 0.987 0.968 0.943 0.978 0.984 0.891
Zimbabwe 0.970 0.952 0.877 0.941 0.972 0.714
Table S6: Benchmark results on the WGND test sets for the CNB and CCT classifiers. Columns AUC, F1, and GAT represent the ROC AUC, F1 score, and gender assignment rate, respectively.

### v2

@@ Acknowledgments
Country AUC F1 $\text{GAR}_{\text{f}}$ $\text{GAR}_{\text{m}}$ AUC F1 $\text{GAR}_{\text{f}}$ $\text{GAR}_{\text{m}}$
Albania 0.984 0.963 0.902 0.911 0.950 0.963 0.853 0.869
Algeria 0.965 0.951 0.880 0.931 0.945 0.953 0.847 0.911
Australia 0.995 0.978 0.964 0.947 0.992 0.994 0.891 0.931
Austria 0.996 0.984 0.962 0.978 0.995 0.996 0.927 0.967
Bahrain 0.998 0.998 0.960 0.956 0.969 0.969 0.824 0.891
Belarus 0.999 0.999 0.951 0.980 0.997 0.998 0.989 0.967
Belgium 0.982 0.971 0.887 0.873 0.979 0.985 0.896 0.864
Bosnia and Herzegovina 0.986 0.974 0.889 0.927 0.979 0.978 0.896 0.920
Bulgaria 0.998 0.993 0.996 0.984 0.987 0.991 0.933 0.798
Canada 0.987 0.973 0.893 0.918 0.984 0.986 0.851 0.918
Cyprus 0.976 0.974 0.749 0.889 0.971 0.970 0.847 0.931
Czechia 0.999 0.995 0.967 0.951 0.999 0.999 0.953 0.909
Denmark 0.989 0.977 0.884 0.931 0.987 0.987 0.864 0.893
Egypt 0.988 0.968 0.924 0.953 0.945 0.944 0.733 0.849
Estonia 0.998 0.994 0.956 0.969 0.991 0.992 0.847 0.851
France 0.995 0.986 0.960 0.916 0.991 0.995 0.902 0.858
Germany 0.995 0.991 0.944 0.933 0.995 0.995 0.931 0.940
Ghana 0.994 0.982 0.938 0.920 0.975 0.986 0.876 0.896
Iceland 0.999 0.997 0.984 0.987 0.988 0.990 0.902 0.918
Iran, Islamic Republic of 0.981 0.974 0.860 0.844 0.933 0.941 0.769 0.856
Iraq 0.841 0.938 0.347 0.364 0.906 0.927 0.618 0.780
Ireland 0.999 0.997 0.958 0.971 0.998 0.998 0.918 0.969
Israel 0.965 0.953 0.822 0.889 0.941 0.952 0.811 0.820
Italy 0.999 0.999 0.998 0.999 0.999 0.999 0.991 0.860
Jamaica 0.978 0.969 0.838 0.833 0.964 0.972 0.762 0.816
Japan 0.981 0.956 0.904 0.964 0.967 0.970 0.733 0.882
Jordan 0.995 0.986 0.971 0.958 0.951 0.954 0.736 0.867
Kenya 0.914 0.895 0.918 0.922 0.907 0.896 0.938 0.929
Kuwait 0.985 0.968 0.898 0.927 0.928 0.942 0.631 0.876
Lebanon 0.997 0.992 0.942 0.976 0.926 0.931 0.802 0.929
Lesotho 0.975 0.965 0.887 0.847 0.687 0.858 0.147 0.213
Lithuania 0.994 0.991 0.998 0.993 0.994 0.994 0.920 0.971
Moldova, Republic of 0.993 0.986 0.987 0.960 0.985 0.986 0.969 0.913
Montenegro 0.983 0.955 0.916 0.920 0.955 0.959 0.864 0.889
Morocco 0.989 0.985 0.898 0.893 0.973 0.976 0.911 0.913
Netherlands 0.981 0.973 0.864 0.931 0.980 0.984 0.864 0.896
New Zealand 0.998 0.992 0.962 0.942 0.989 0.991 0.898 0.893
Nigeria 0.954 0.923 0.862 0.869 0.933 0.952 0.604 0.742
Norway 0.991 0.980 0.980 0.987 0.984 0.986 0.893 0.887
Oman 0.999 0.997 0.958 0.980 0.908 0.918 0.742 0.851
Papua New Guinea 0.983 0.969 0.887 0.907 0.946 0.970 0.764 0.784
Philippines 0.981 0.978 0.873 0.849 0.958 0.968 0.784 0.822
Poland 0.992 0.990 0.982 0.973 0.991 0.993 0.973 0.929
Portugal 0.996 0.992 0.933 0.913 0.992 0.994 0.924 0.924
Romania 0.980 0.967 0.922 0.920 0.962 0.973 0.851 0.787
Russian Federation 0.999 0.996 0.993 0.998 0.998 0.999 0.964 0.929
Saudi Arabia 0.981 0.975 0.858 0.933 0.878 0.901 0.684 0.849
Serbia 0.997 0.994 0.967 0.987 0.984 0.979 0.949 0.938
Somalia 0.824 0.883 0.280 0.480 0.964 0.971 0.749 0.867
South Africa 0.979 0.962 0.884 0.858 0.919 0.966 0.624 0.671
Spain 0.990 0.978 0.944 0.940 0.986 0.984 0.920 0.924
Sri Lanka 0.884 0.927 0.500 0.473 0.858 0.930 0.562 0.582
Sweden 0.996 0.988 0.922 0.949 0.988 0.992 0.933 0.909
Switzerland 0.999 0.992 0.989 0.991 0.999 0.999 0.962 0.933
Syrian Arab Republic 0.917 0.888 0.822 0.840 0.943 0.954 0.653 0.853
Tunisia 0.965 0.953 0.898 0.844 0.964 0.966 0.807 0.871
Türkiye 0.990 0.975 0.924 0.936 0.972 0.977 0.811 0.898
United Arab Emirates 0.995 0.991 0.933 0.969 0.942 0.950 0.742 0.902
United Kingdom 0.998 0.991 0.967 0.964 0.996 0.998 0.876 0.918
United States 0.987 0.968 0.942 0.944 0.978 0.984 0.849 0.933
Zimbabwe 0.970 0.952 0.891 0.862 0.941 0.972 0.771 0.658
Table S6: Benchmark results on the WGND test sets for the CNB and CCT classifiers. Columns AUC and F1 represent the ROC AUC and F1 score, respectively. $\text{GAR}_{\text{f}}$ and $\text{GAR}_{\text{m}}$ represent the gender assignment rates for the female and male test samples, respectively.